OpenStudy (maxdrizner):

Which of the following best defines 2/2^3

OpenStudy (solomonzelman):

\[\frac{2}{2^3},~~~->~~~\frac{2}{8},~~~->~~~\frac{1}{4}.\]

OpenStudy (maxdrizner):

these are the answers Cube root of 4 Square root of 4 Cube root of 2 Square root of 2

OpenStudy (solomonzelman):

NONE. Perhaps your question is different from what you actually wrote?

OpenStudy (maxdrizner):

its wanting me to define 2 over 2 to the 3rd power

OpenStudy (solomonzelman):

\[(\frac{2}{2})^3,~~~->~~~~(1)^3~~~->~~~~1.\]

OpenStudy (maxdrizner):

these are the answers Cube root of 4 Square root of 4 Cube root of 2 Square root of 2

OpenStudy (solomonzelman):

Well, again, NONE of these are the answers to the\[(\frac{2}{2})^3\]

OpenStudy (maxdrizner):

well i dont have a none of the above answers

OpenStudy (solomonzelman):

I know, but\[(\frac{2}{2})^3=1\]Anybody will tell you the same thing.

OpenStudy (solomonzelman):

Your school must be HORRIBLE....

OpenStudy (maxdrizner):

well its an online class

OpenStudy (solomonzelman):

Oh, I don't have the username for that, I know you can sopy paste the link, and it will open right away, because you're registered, but I am not.... I had the same with connexus once.

OpenStudy (maxdrizner):

darn

OpenStudy (maxdrizner):

maybe i can copy in in note pad

OpenStudy (solomonzelman):

You can put it in equations, in a reply.

OpenStudy (maxdrizner):

\[\frac{ 2 }{ 2^3 }\]

OpenStudy (solomonzelman):

, then, as a showed before, it is 1/4.

OpenStudy (maxdrizner):

ok

OpenStudy (solomonzelman):

Look at your choices, maybe they are written incorrectly?

OpenStudy (maxdrizner):

these are my choices Cube root of 4 Square root of 4 Cube root of 2 Square root of 2

OpenStudy (solomonzelman):

\[AGAIN,~~~~~~(\frac{2}{2})^3=1~~~\And~~~\frac{2}{2^3}=\frac{1}{4},~~~So~~all~~of~~the~~choices~~are~~wrong.\]

OpenStudy (maxdrizner):

STAWP YERING ART MEH

OpenStudy (solomonzelman):

Id get this reply, Sorry.